# Polarized semi-inclusive deep-inelastic scattering at NNLO in QCD

### v2

@@ II Kinematics of polarized SIDIS
×$\times\,D_{p^{\prime}}^{h}\left(\frac{z}{\hat{z}},\mu_{A}^{2}\right)\Delta\mathcal{C}_{p^{\prime}p}\left(\hat{x},\hat{z},Q^{2},\mu_{R}^{2},\mu_{F}^{2},\mu_{A}^{2}\right).$ (4)
@@
×$\times\,D_{p^{\prime}}^{h}\left(\frac{z}{\hat{z}},\mu_{A}^{2}\right)\mathcal{C}^{T}_{p^{\prime}p}\left(\hat{x},\hat{z},Q^{2},\mu_{R}^{2},\mu_{F}^{2},\mu_{A}^{2}\right).$ (5)
@@ IV Results
Measurements of SIDIS for various species of final state hadrons on proton targets were performed by CERN COMPASS [18] and the DESY HERMES [19] fixed-target experiments. The experiments present their results in terms of the longitudinal double spin asymmetry $A_{1}^{h}$ (3). COMPASS [18] measured at $\sqrt{s}=17.4$ GeV and included all hadrons produced in the range $0.850.2<z<$ 0.85. The HERMES [19] data were taken at $\sqrt{s}=7.25$ GeV and included hadrons in the range $0.850.2<z<$ 0.85. The asymmetry $A_{1}^{h}$ is evaluated at central bin values in $x$ and $Q^{2}$, justified by a relatively narrow range in $Q^{2}$ for each bin due to the kinematical restrictions of the experiments, at $\mu_{R}=\mu_{F}=\mu_{A}=\sqrt{Q^{2}}$.
@@
To study the perturbative convergence of the polarized SIDIS predictions at higher energies, Fig. 3 displays $g_{1}^{\pi^{+}}$ (integrated over $0.850.2<z<$ 0.85) of the proton for EIC-like kinematics with $\sqrt{s}=45$ GeV. We consider several values of $Q^{2}$, and for each value of $Q^{2}$ we plot predictions for values of $x$ constrained by the requirement $0.90.5<y<$ 0.9. Using the polarized PDFs and FFs at NNLO throughout allows us to assess the uncertainty on the theory predictions in a consistent manner through a seven-point scale variation on $\mu_{R}$ and $\mu_{F}=\mu_{A}$ around their central value $\sqrt{Q^{2}}$, discarding opposite variations of any pair of scales.
@@
We observe very good perturbative convergence of the predictions. While the NLO contributions enhance the LO prediction by 5–30%, we note that in the range $0.50.03<x<$ 0.5, inclusion of the NNLO contribution changes the predictions by less than ±$\pm 5\%$, slightly increasing towards smaller or larger values of $x$. The scale uncertainty at NLO ranges between ±$\pm 5\%$ at low $x$ values to ±$\pm 15\%$ at high $x$. At NNLO, it is reduced throughout the kinematical range to ±$\pm 2\%$ at low $(x,Q^{2})$ and at most ±$\pm 8\%$ at larger values. The non-trivial kinematical shape of the NNLO corrections highlights their importance for precision physics studies at the EIC.

### v3

@@ II Kinematics of polarized SIDIS
$\times\,D_{p^{\prime}}^{h}\left(\frac{z}{\hat{z}},\mu_{A}^{2}\right)\Delta\mathcal{C}_{p^{\prime}p}\left(\hat{x},\hat{z},Q^{2},\mu_{R}^{2},\mu_{F}^{2},\mu_{A}^{2}\right).$ (4)
@@
$\times\,D_{p^{\prime}}^{h}\left(\frac{z}{\hat{z}},\mu_{A}^{2}\right)\mathcal{C}^{T}_{p^{\prime}p}\left(\hat{x},\hat{z},Q^{2},\mu_{R}^{2},\mu_{F}^{2},\mu_{A}^{2}\right).$ (5)
@@ IV Results
Measurements of SIDIS for various species of final state hadrons on proton targets were performed by CERN COMPASS [18] and the DESY HERMES [19] fixed-target experiments. The experiments present their results in terms of the longitudinal double spin asymmetry $A_{1}^{h}$ (3). COMPASS [18] measured at $\sqrt{s}=17.4$ GeV and included all hadrons produced in the range $0.2<z<0.85$. The HERMES [19] data were taken at $\sqrt{s}=7.25$ GeV and included hadrons in the range $0.2<z<0.85$. The asymmetry $A_{1}^{h}$ is evaluated at central bin values in $x$ and $Q^{2}$, justified by a relatively narrow range in $Q^{2}$ for each bin due to the kinematical restrictions of the experiments, at $\mu_{R}=\mu_{F}=\mu_{A}=\sqrt{Q^{2}}$.
@@
To study the perturbative convergence of the polarized SIDIS predictions at higher energies, Fig. 3 displays $g_{1}^{\pi^{+}}$ (integrated over $0.2<z<0.85$) of the proton for EIC-like kinematics with $\sqrt{s}=45$ GeV. We consider several values of $Q^{2}$, and for each value of $Q^{2}$ we plot predictions for values of $x$ constrained by the requirement $0.5<y<0.9$. Using the polarized PDFs and FFs at NNLO throughout allows us to assess the uncertainty on the theory predictions in a consistent manner through a seven-point scale variation on $\mu_{R}$ and $\mu_{F}=\mu_{A}$ around their central value $\sqrt{Q^{2}}$, discarding opposite variations of any pair of scales.
@@
We observe very good perturbative convergence of the predictions. While the NLO contributions enhance the LO prediction by 5–30%, we note that in the range $0.03<x<0.5$, inclusion of the NNLO contribution changes the predictions by less than $\pm 5\%$, slightly increasing towards smaller or larger values of $x$. The scale uncertainty at NLO ranges between $\pm 5\%$ at low $x$ values to $\pm 15\%$ at high $x$. At NNLO, it is reduced throughout the kinematical range to $\pm 2\%$ at low $(x,Q^{2})$ and at most $\pm 8\%$ at larger values. The non-trivial kinematical shape of the NNLO corrections highlights their importance for precision physics studies at the EIC.